# SIR-RL: Reinforcement Learning for Optimized Policy Control during Epidemiological Outbreaks in Emerging Market and Developing Economies

### v1

@@ 5.1 Dataset
The population-level epidemiological data can be obtained from the “Our World In Data COVID-19” dataset: https://ourworldindata.org/coronavirus or more specifically: https://github.com/owid/covid-19-data/blob/master/public/data/owid-covid-data.csv [67]. Data for the total cases, and recovered was acquired by scraping the Worldometers website [31] used Internet Archive [73]. Quaterly GDP data can be obtained from the “Organisation for Economic Co-operation and Development”: https://www.oecd-ilibrary.org/economics/data/main-economic-indicators/main-economic-indicators-complete-database_data-00052-en [74].

### v2

@@ 5.1 Dataset
The population-level epidemiological data can be obtained from the “Our World In Data COVID-19” dataset: https://ourworldindata.org/coronavirus or more specifically: https://github.com/owid/covid-19-data/blob/master/public/data/owid-covid-data.csv [67]. Data for the total cases, and recovered was acquired by scraping the Worldometers website [31] using Internet Archive [73]. Quaterly GDP data can be obtained from the “Organisation for Economic Co-operation and Development”: https://www.oecd-ilibrary.org/economics/data/main-economic-indicators/main-economic-indicators-complete-database_data-00052-en [74].
@@ 5.2 Code
5.3 Data Availablity
All the data used in the manuscript has been obtained from open data sources and has been cited in the manuscript itself.